Consider a firm with the following short-run costs:

Quantity

Variable Cost

Total Cost

1

30

90

2

50

110

3

90

150

4

140

200

5

200

260

Calculate the firm's profit/loss, if the market price is $50 (MR=$50).
Answer #1

Profit is the difference between the total revenue and the total cost of the production. The total cost is given in the table and we want the total revenue.

Profit=TR-TC.

The firm naturally produce where the marginal revenue equals the marginal cost of the production. This is the profit maximization condition for a firm. So here the marginal revenue is same as the price of the product, and the marginal cost is the addition made to the total cost when an additional unit of the commodity is produced. The marginal cost is calculated given below.

TC-90-110-150 200-360 1|2|3 4 5TR=P\times Q

50 × 4

200

PROFIT=TR-TC

200-200

  =0.

Ans: There is zero economic profit.